OpenStudy (anonymous):

Verify that quadrilateral BCDE is a rhombus with vertices B(5, -1), C(-2, 0), D( -1, 7) and E(6, 6) by showing that all four sides are equal.

OpenStudy (anonymous):

Show that all sides are equal and its diagonals are perpendicular to each other.

OpenStudy (anonymous):

Calculate the lengths of all sides by using distance formula d=sqrt((x2-x1)^2+(y2-y1)^2) BC=sqrt((-2-5)^2+(0-(-1))^2)=sqrt(49+1)=sqrt(50)=5sqrt(2) CD=((-1-(-2))^2+(7-0)^2)=sqrt(1+49)=sqrt(50)=5sqrt(2) and so on ...

OpenStudy (anonymous):

so wouldnt it be 11?